OpenStudy (anonymous):

Solve for the variable. -x/9=5/3

OpenStudy (anonymous):

x=-15..

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

what about this one The volume of a cylinder is V = pie r ^2 h. Find h if pie =3.14, V = 128, and r = 8 Round to the nearest hundredth.

OpenStudy (anonymous):

h=128/(3.14)(64)=.63

OpenStudy (anonymous):

here are the choices. A. 0.56 B. 0.64 C. 0.58 D. 0.52

OpenStudy (anonymous):

So B.

OpenStudy (anonymous):

.64..

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

is this right \[\frac{ x }{ 10 } + \frac{ x }{ 14 }\]

OpenStudy (anonymous):

for this question Crane A can unload a ship in 10 hours and crane B can unload it in 14 h. How long will it take the two cranes to unload the ship working together? What is the correct equation to solve this problem?